OpenStudy (anonymous):

The couch they like costs $2700. The salesperson told the Johnsons the couch will be on sale for 20% off. the sales tax rate in their city is 7.5%, how much will the Johnsons save including sales tax?

OpenStudy (anonymous):

2700*0.075=202.5 2700+202.5=2902.5 2902.5*.20=580.5 2902.5-580.5=2322

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

or 1998